Two sides of a garden bed are to be lined with logs. Three logs are needed for each side but 1m is cut off the third log to make sure that the logs don't extend beyond the garden. The width of the garden is 3m and its area is 42mē. How long is each log?
Answer by josgarithmetic(39617)

Garden dimensions are 3 meters by 42/3=21*2/3=14 meters. The rest is not clear; which two sides Opposite or adjascent? Total log length for all six logs?